10. A 1500 MS2 resistor is placed across the terminals of a battery, and a terminal voltage of 2.50 (1 point) V is measured. when this resistor is replaced with a 5.00 Ω resistor, the terminal voltage decreases to 1.75 V Find the terminal voltage when a 7.00 Ω resistor is used. O2.50 V 01.91 V 01.67 V O2.14 V
Answer #1

1.91 V is the correct answer to the given problem.

Let V_i and r be the internal EMF and resistance of the battery.

For the 1500 M Omega resistance (say R_1), the equation for internal EMF of battery is,

V_i = I_1R_1 + I_1r

where I_1 is the current flowing in the circuit with R_1 resistance connected.

The current flowing in the circuit is given as,

I_1 = rac{V_1}{R_1}

with V_1 as the terminal voltage in this case.

I_1 = rac{2.5}{1500 imes10^6}\ = 1.67 imes 10^{-9} A

Similarly, for the 5 Omega resistance (say R_2), the equation for internal EMF of battery is,

V_i = I_2R_2 + I_2r

I_2 = rac{V_2}{R_2}

= rac{1.75}{5}\ = 0.35 A

Equating both the equations to find r,

I_1R_1 + I_1r = I_2R_2 + I_2r\ r = rac{I_1R_1- I_2R_2}{I_2 - I_1}

=rac{2.5 - 1.75}{0.35 - 1.67 imes 10^{-9}}\ = 2.14 Omega

The internal EMF of the battery may be obtained from the second equation for V_i as,

Vi = 0.35 * 5 2,5 У 0.35 * 2.14

In the case of 7.00 Omega resistor (say R) connected across the battery, the current (I) in the circuit is given as,

I = rac{V_i}{R + r}\ = rac{2.5}{7.00+2.14}\ = 0.27 A

The terminal voltage (V) in this case is,

0.27*7 1.9 V